# Archetypal SAE: Adaptive and Stable Dictionary Learning for Concept Extraction in Large Vision Models

### v1

@@ Abstract
Matthew Kowal FAR AI, York University Victor Boutin CerCo CNRS Isabel Papadimitriou Kempner Institute Harvard University Binxu Wang Kempner Institute Harvard University
Martin Wattenberg Harvard University, Google DeepMind Demba Ba Kempner Institute Harvard University Talia Konkle Dept. of Psychology & Kempner Institute Harvard University
footnotemark: footnotetext: These authors contributed equally to this work.footnotemark: footnotetext: work done at Harvard.footnotemark: footnotetext:
 Preprint. Under review.
@@ 4 Towards Archetypal SAEs
In contrast to above, the Archetypal dictionary imposes a crucial geometric restriction: every dictionary atom is constrained to lie within the convex hull of $\bm{A}$, i.e., each dictionary atom $\bm{D}_{i}$ is a convex combination of samples’ representations. Thus, once multiplied by a nonnegative $\bm{Z}$, the reconstructions $\bm{Z}\bm{D}$ remain within the conic hull of $\bm{A}$. This anchoring within the data manifold precludes the emergence of pathological or out‐of‐sample directions, yielding stability gains shown empirically in Figure 3 and formalized in Proposition E.2. Concretely, we always have:
@@
Moreover, one can restrict $\bm{D}$ to be formed from a subset $\bm{C}\subseteq\mathbb{R}^{n^{\prime}\times d}$ (rather than all of $\bm{A}$) without losing expressivity, provided $\bm{C}$ contains the extreme points of $\bm{A}$ (Dubins, 1962). Indeed, in that case $\mathrm{cone}(\bm{C})=\mathrm{cone}(\bm{A})$, ensuring the same representational power (see Proposition E.1).
@@
To maintain the desirable properties of A-SAE while addressing scalability, we propose using a reduced subset of points $\bm{C}$, with $n^{\prime}\ll n$, chosen as centroids of $\bm{A}$ obtained via K-Means. We fix $\bm{C}$ and train only $\bm{W}$. We apply a ReLU activation and normalize each row of $\bm{W}$ at each step to ensure that $\bm{W}\in\Omega_{k,n}$. Experiments indicate that K-Means forms the most reliable method for distilling $\bm{A}$ into $\bm{C}$ (see Figure 11), compared to alternatives such as isolation forests (Liu et al., 2008), convex hull computation in reduced dimensions, or outlier detection methods (Scholkopf et al., 1999; Breunig et al., 2000) (details in Appendix C). Hereafter, we use A-SAE to refer to this implementation where we directly optimize $\bm{W}$ to find the best convex combination of points that reconstruct data.
@@ 5 Experiments
We evaluate five models: DINOv2 (Darcet et al., 2023; Oquab et al., 2023), SigLip (Zhai et al., 2023), ViT (Dosovitskiy et al., 2020), ConvNeXt (Liu et al., 2022), and ResNet50 (He et al., 2016), sourced from the timm library (Wightman et al., 2019). Unless specified, we trained overcomplete dictionaries with size $k=5\times$ the feature dimension (e.g., $768\times 5$ for DINOv2 and $2048\times 5$ for ConvNeXt). Models were trained on the ImageNet dataset ($\sim 1.28$M images), resulting in over $60$M tokens per epoch for ConvNeXt ($7\times 7$ tokens/image) and over $250$M tokens per epoch for DINOv2 ($14\times 14$ patches/image) across 50 epochs. Semi-NMF and Convex-NMF were trained using gradient descent with accumulation and $\ell_{1}$ regularization to control sparsity, while the RA-SAE was applied atop a TopK SAE to maintain consistent sparsity. To compute $\bm{C}$, we applied K-Means clustering to the entire dataset, reducing it to $32,\!000$ centroids, which achieved reconstruction error comparable to the unconstrained SAE. The data matrix $\bm{A}$ was element-wise standardized.
@@ 5.1 Dictionary Learning Metrics
To improve SAEs, it is essential to deepen our understanding of the solutions they yield. We thus evaluate both standard and novel sets of metrics that evaluate dictionary learning methods across four key dimensions: (i) sparse reconstruction, (ii) consistency, (iii) structure in the dictionary ($\bm{D}$), and (iv) structure in the codes ($\bm{Z}$), all reported in Tab. 1 (see App. B for formal definitions). Overall, we find Archetypal SAEs achieve a strong balance between reconstruction performance, consistency, and identification of structure.
@@
ii) Consistency. Excelling in reconstruction does not guarantee that the learned solution aligns with the underlying data distribution or is stable across training runs. To this end, we measure stability, which assesses the consistency of learned dictionaries across different initializations or perturbations in the data, and the OOD score, which quantifies how close the dictionary atoms are to real data points, i.e., whether learned concepts remain grounded and interpretable. Our findings indicate that SAEs perform poorly in consistency, showing both low stability (as evidenced in Figure 3) and suboptimal OOD scores. In contrast, Archetypal SAEs significantly enhance stability and OOD score without sacrificing reconstruction performance. In Proposition E.4, we provide theoretical arguments showing that, under mild assumptions, the OOD score of Archetypal SAEs is inherently lower-bounded, ensuring that learned dictionary atoms remain well-anchored within the data.
iii) Structure in the Dictionary ($\bm{D}$). This dimension examines whether the learned dictionary exhibits meaningful patterns. A well-structured dictionary may reveal meta-concepts or decomposable higher-level abstractions. Metrics such as stable rank, effective rank, and coherence provide insights into the compactness and interpretability of the dictionary. In particular, among solutions with comparable reconstruction performance, a more structured dictionary is preferable due to its potential for higher interpretability and organization. Across runs, methods like CNMF, SNMF, and Archetypal dictionary learning consistently yield dictionaries with better structure, reflecting their capacity for capturing higher-level patterns within the data. We also provide theoretical arguments in Proposition E.3, demonstrating that the rank of Archetypal dictionaries is inherently bounded by the rank of the data matrix.
@@ 5.3 Soft Identifiability Benchmark
Motivated by the above, we next propose a Soft Identifiability Benchmark that assesses the capability of SAEs to disentangle synthetic image mixtures with known underlying concepts. Specifically, we construct twelve synthetic datasets, each comprising images formed by collaging four distinct objects (e.g., different types of gems) sampled from a pre-defined set. Each dataset is processed through a pre-trained vision model to obtain pooled activations, which serve as inputs for the SAE. Ideally, when trained on these activations, the SAE is able to recover the original objects as distinct concepts within its dictionary. To evaluate the same, we assess performance by checking whether each object class has a corresponding concept that activates above a threshold $\lambda$ when an object $y_{j}$ appears in the image. Formally, for each image, we feed it into a model and then into the trained SAE to get a concept-label pair $(\bm{z},\bm{y})$, where $\bm{z}\in\mathbb{R}^{k}$ represents the $k$ concept values and $\bm{y}\in\mathbb{R}^{c}$ denotes the $c$ class labels. We then define the accuracy for the class $j$ as: $\text{Accuracy}_{j}=\max_{\lambda\in\mathbb{R},i\in[k]}~{}~{}\mathbb{P}_{(\bm{z},\bm{y})}((z_{i}>\lambda)=y_{j})$.
Figure 7: Exotic Concepts in DinoV2, among the $16,000$ concepts found by RA-SAE, A) Highlights tokens in shadows of dogs, suggesting shadow-based features, potentially used for depth reasoning. B) A “barber” concept exclusively active for barbers, not their clients. C) A fine-grained visual concept focusing on petal edges or contours.
@@
Table 3: Soft Identifiability benchmark Across Models and Methods. This table presents the average accuracy scores for various methods evaluated across four different models: DINO, ResNet, SigLIP, and ViT. Best-performing scores for each model are in bold and second best are underlined. Full results are available in Appendix G.
Intuitively, the evaluation above gauges whether each object is disentangled into its “own box”, in that a threshold $\lambda$ enables one-hot classification for all objects. In each of the 12 synthetic datasets, the size of the dictionary is set to the exact number of distinct objects (from 9 to 20 objects) and we generate datasets of 2,000 samples. An example of the dataset and the procedure is given in Figure 6.
Results are shown in Table 3. Classical dictionary learning methods and standard SAEs exhibit comparable performance, struggling to accurately recover the true generative factors. In contrast, RA-SAE consistently achieves higher accuracy in recovering the underlying object classes across all synthetic datasets. This indicates that RA-SAE not only stabilizes the dictionary, but also enhances its ability to discern the true underlying structure of the data, thereby improving the disentanglement of mixed objects.
@@
5.4 Qualitative Results
Beyond quantitative evaluations on synthetic datasets, we also provide qualitative results to illustrate the meaningfulness of the concepts learned by RA-SAE on DinoV2 with natural images. RA-SAE uncovers unexpected concepts, such as shadow-based features (potentially linked to depth reasoning), a context-dependent “barber” concept (activating for barbers but not clients), and fine-grained edge detection in flower petals (Figure 7). It also learns more structured within-class distinctions (e.g., separating rabbit ears, faces, and paws) compared to TopK-SAEs (Figure 8). Finally, its dictionary forms clear clusters, grouping semantically related features like fine-grained animal faces or spatial concepts such as “underneath” (Figure 10).
@@ Appendix A Qualitative Examples
This section provides qualitative insights into the concepts learned by DinoV2 (Oquab et al., 2023; Darcet et al., 2023). We trained an Archetypal SAE on DinoV2-B with 4 registers and reported key performance metrics in Figure 9.
In addition, we identified surprising and specific concepts among the $16,000$ dictionary atoms, shown in Figure LABEL:fig:surprising. These include: A) A concept highlighting tokens corresponding to shadows of dogs, suggesting that DinoV2 may use shadows as a feature, potentially contributing to its depth estimation or 3D reasoning capabilities. B) A “barber” concept that activates exclusively on tokens representing barbers but not on the individuals receiving haircuts or shaves. C) A fine-grained visual concept that activates along the contours or edges of flower petals.
Appendix B Formal Definitions of Metrics
@@ B.3 Structure in the Dictionary ($\bm{D}$)
Appendix C Distilling $\bm{A}$ into $\bm{C}$
@@
Appendix D Implementation Details
@@
Appendix E Theoretical properties of Archetypal Dictionary
@@
E.1 Geometric interpretation of A-SAE
Proposition E.1 (Archetypal Dictionary, Convex and Conic Hulls).
@@ E.1 Geometric interpretation of A-SAE
E.2 Stability of Archetypal Dictionary
Proposition E.2 (Geometric Stability of Archetypal Dictionaries).
@@ E.2 Stability of Archetypal Dictionary
E.3 Controlled Rank and Stability in A-SAE Dictionaries

### v2

@@ Abstract
Matthew Kowal FAR AI York University Victor Boutin CerCo CNRS Isabel Papadimitriou Kempner Institute Harvard University Binxu Wang Kempner Institute Harvard University
Martin Wattenberg Harvard University Google DeepMind Demba Ba Kempner Institute Harvard University Talia Konkle Dept. of Psychology & Kempner Institute Harvard University
footnotemark: footnotetext: These authors contributed equally to this work.footnotemark: footnotetext: work done at Harvard.footnotemark: footnotetext: Proceedings of the $\mathit{42}^{nd}$ International Conference on Machine Learning, Vancouver, Canada. PMLR 267, 2025. Copyright 2025 by the author(s).
@@ 4 Towards Archetypal SAEs
In contrast to above, the Archetypal dictionary imposes a crucial geometric restriction: every dictionary atom is constrained to lie within the convex hull of $\bm{A}$, i.e., each dictionary atom $\bm{D}_{i}$ is a convex combination of samples’ representations. Thus, once multiplied by a nonnegative $\bm{Z}$, the reconstructions $\bm{Z}\bm{D}$ remain within the conic hull of $\bm{A}$. This anchoring within the data manifold precludes the emergence of pathological or out‐of‐sample directions, yielding stability gains shown empirically in Figure 3 and formalized in Proposition F.2. Concretely, we always have:
@@
Moreover, one can restrict $\bm{D}$ to be formed from a subset $\bm{C}\subseteq\mathbb{R}^{n^{\prime}\times d}$ (rather than all of $\bm{A}$) without losing expressivity, provided $\bm{C}$ contains the extreme points of $\bm{A}$ (Dubins, 1962). Indeed, in that case $\mathrm{cone}(\bm{C})=\mathrm{cone}(\bm{A})$, ensuring the same representational power (see Proposition F.1).
@@
To maintain the desirable properties of A-SAE while addressing scalability, we propose using a reduced subset of points $\bm{C}$, with $n^{\prime}\ll n$, chosen as centroids of $\bm{A}$ obtained via K-Means. We fix $\bm{C}$ and train only $\bm{W}$. We apply a ReLU activation and normalize each row of $\bm{W}$ at each step to ensure that $\bm{W}\in\Omega_{k,n}$. Experiments indicate that K-Means forms the most reliable method for distilling $\bm{A}$ into $\bm{C}$ (see Figure 11), compared to alternatives such as isolation forests (Liu et al., 2008), convex hull computation in reduced dimensions, or outlier detection methods (Scholkopf et al., 1999; Breunig et al., 2000) (details in Appendix D). Hereafter, we use A-SAE to refer to this implementation where we directly optimize $\bm{W}$ to find the best convex combination of points that reconstruct data.
@@ 5 Experiments
We evaluate five models: DINOv2 (Darcet et al., 2023; Oquab et al., 2023), SigLip (Zhai et al., 2023), ViT (Dosovitskiy et al., 2020), ConvNeXt (Liu et al., 2022), and ResNet50 (He et al., 2016), sourced from the timm library (Wightman et al., 2019). Unless specified, we trained overcomplete dictionaries with size $k=5\times$ the feature dimension (e.g., $768\times 5$ for DINOv2 and $2048\times 5$ for ConvNeXt). Models were trained on the ImageNet dataset ($\sim 1.28$M images), resulting in over $60$M tokens per epoch for ConvNeXt ($7\times 7$ tokens/image) and over $250$M tokens per epoch for DINOv2 ($14\times 14$ patches/image) across 50 epochs. Semi-NMF and Convex-NMF were trained using gradient descent with accumulation and $\ell_{1}$ regularization to control sparsity, while the RA-SAE was applied atop a TopK SAE to maintain consistent sparsity. To compute $\bm{C}$, we applied K-Means clustering to the entire dataset, reducing it to $32,\!000$ centroids, which achieved reconstruction error comparable to the unconstrained SAE. The data matrix $\bm{A}$ was element-wise standardized. All the SAEs were trained using the Overcomplete library [^6] https://github.com/KempnerInstitute/overcomplete.
@@ 5.1 Dictionary Learning Metrics
To improve SAEs, it is essential to deepen our understanding of the solutions they yield. We thus evaluate both standard and novel sets of metrics that evaluate dictionary learning methods across four key dimensions: (i) sparse reconstruction, (ii) consistency, (iii) structure in the dictionary ($\bm{D}$), and (iv) structure in the codes ($\bm{Z}$), all reported in Tab. 1 (see App. C for formal definitions). Overall, we find Archetypal SAEs achieve a strong balance between reconstruction performance, consistency, and identification of structure.
@@
ii) Consistency. Excelling in reconstruction does not guarantee that the learned solution aligns with the underlying data distribution or is stable across training runs. To this end, we measure stability, which assesses the consistency of learned dictionaries across different initializations or perturbations in the data, and the OOD score, which quantifies how close the dictionary atoms are to real data points, i.e., whether learned concepts remain grounded and interpretable. Our findings indicate that SAEs perform poorly in consistency, showing both low stability (as evidenced in Figure 3) and suboptimal OOD scores. In contrast, Archetypal SAEs significantly enhance stability and OOD score without sacrificing reconstruction performance. In Proposition F.4, we provide theoretical arguments showing that, under mild assumptions, the OOD score of Archetypal SAEs is inherently lower-bounded, ensuring that learned dictionary atoms remain well-anchored within the data.
iii) Structure in the Dictionary ($\bm{D}$). This dimension examines whether the learned dictionary exhibits meaningful patterns. A well-structured dictionary may reveal meta-concepts or decomposable higher-level abstractions. Metrics such as stable rank, effective rank, and coherence provide insights into the compactness and interpretability of the dictionary. In particular, among solutions with comparable reconstruction performance, a more structured dictionary is preferable due to its potential for higher interpretability and organization. Across runs, methods like CNMF, SNMF, and Archetypal dictionary learning consistently yield dictionaries with better structure, reflecting their capacity for capturing higher-level patterns within the data. We also provide theoretical arguments in Proposition F.3, demonstrating that the rank of Archetypal dictionaries is inherently bounded by the rank of the data matrix.
@@ 5.3 Soft Identifiability Benchmark
Table 3: Soft Identifiability benchmark Across Models and Methods. This table presents the average accuracy scores for various methods evaluated across four different models: DINO, ResNet, SigLIP, and ViT. Best-performing scores for each model are in bold and second best are underlined. Full results are available in Appendix H.
Motivated by the above, we propose a Soft Identifiability Benchmark. Specifically, we construct twelve synthetic datasets, each comprising images formed by collaging four distinct objects (e.g., different types of gems) sampled from a pre-defined set. Each dataset is processed through a pre-trained vision model to obtain pooled activations. Ideally, when trained on these activations, the SAE is able to recover the original objects as distinct concepts within its dictionary. We then assess performance by checking whether each object class has a corresponding concept that activates above a threshold $\lambda$ when an object $y_{j}$ appears in the image. Formally, for each image, we feed it into a model and then into the trained SAE to get a concept-label pair $(\bm{z},\bm{y})$, where $\bm{z}\in\mathbb{R}^{k}$ represents the $k$ concept values and $\bm{y}\in\mathbb{R}^{c}$ denotes the $c$ class labels. We then define the accuracy for the class $j$ as: $\text{Accuracy}_{j}=\max_{\lambda\in\mathbb{R},i\in[k]}~{}~{}\mathbb{P}_{(\bm{z},\bm{y})}((z_{i}>\lambda)=y_{j})$.
5.4 Qualitative Examples
@@
This section provides qualitative insights into the concepts learned by DinoV2 (Oquab et al., 2023; Darcet et al., 2023). We trained an Archetypal SAE on DinoV2-B with 4 registers and reported key performance metrics in Figure 10. RA-SAE uncovers unexpected concepts, such as shadow-based features (potentially linked to depth reasoning), a context-dependent “barber” concept (activating for barbers but not clients), and fine-grained edge detection in flower petals (Figure 9). It also learns more structured within-class distinctions (e.g., separating rabbit ears, faces, and paws) compared to TopK-SAEs (Figure 8). Finally, its dictionary forms clear clusters, grouping semantically related features like fine-grained animal faces or spatial concepts such as “underneath” (Figure 7).
In addition, we identified surprising and specific concepts among the $16,000$ dictionary atoms, shown in Figure LABEL:fig:surprising_main. These include: A) A concept highlighting tokens corresponding to shadows of dogs, suggesting that DinoV2 may use shadows as a feature, potentially contributing to its depth estimation or 3D reasoning capabilities. B) A “barber” concept that activates exclusively on tokens representing barbers but not on the individuals receiving haircuts or shaves. C) A fine-grained visual concept that activates along the contours or edges of flower petals.
Figure 9: Exotic Concepts in DinoV2. A) Highlights tokens in shadows of dogs, suggesting shadow-based features and potential depth reasoning. B) A “barber” concept exclusively active for barbers, not their clients. C) A fine-grained concept focusing on petal edges or contours.
@@ 6 Conclusion
Impact Statement
This work confronts a critical obstacle in sparse interpretability methods: instability. Archetypal SAEs enforce geometric constraints that yield dictionaries more consistent across runs, more grounded in the data manifold, and better aligned with true classification and generative directions. These advances enhance both the reliability and scientific utility of sparse autoencoders, enabling stable concept discovery at scale in large vision models. Our benchmarks provide a principled framework for evaluating concept plausibility and identifiability, supporting reproducible and trustworthy progress in interpretability research.
@@ Appendix A Qualitative Examples
Appendix B Extended Related Work on Archetypal Analysis
Archetypal analysis (AA) was introduced by Cutler and Breiman in 1994 as a method to represent observations as convex combinations of extremal “pure types” called archetypes (Cutler & Breiman, 1994). Each archetype lies on the boundary of the data’s convex hull and is itself constrained to be a convex combination of data points, yielding an interpretable factorization where data are explained in terms of extreme exemplars. This approach was proposed as an alternative to principal component analysis for uncovering latent structure, providing data-like representative factors instead of orthogonal directions (Cutler & Stone, 1997). Shortly after its introduction, AA was extended to spatio-temporal settings by identifying archetypes that vary over time (Stone & Cutler, 1996) and (Prabhakaran et al., 2012) proposed an automatic model selection criterion for AA, alleviating the need to fix $k$ a priori. After its initial formulation, AA did not gain widespread popularity, partly due to computational limitations and the availability of more established techniques like $k$-means or NMF. Nonetheless, several works in the 2000s applied AA successfully in scientific domains. For instance (Chan et al., 2003) applied AA to astronomical spectra. In genetics,  (Gimbernat-Mayol et al., 2022) found that archetypes learned from human genotype data corresponded to ancestral population prototypes. AA was also utilized in socio-economic and bibliometric analyses (Seiler & Wohlrabe, 2013), and to model player behavior in video games (Sifa & Bauckhage, 2013). Various enhancements to the basic AA model have been proposed, (Eugster & Leisch, 2011) introduced weighted and robust AA, (Seth & Eugster, 2016) proposed to treats archetypes as latent factors in a probabilistic generative model. In a similar vein, contemporary work has developed AA for binary data specifically (Wedenborg & Mørup, 2025) and  (Epifanio et al., 2019) tackled the presence of missing values. Likewise (Moliner & Epifanio, 2019) formulated robust AA for multivariate functional data.
Another important direction has been the development of non-linear and kernel variants of AA (Mørup & Hansen, 2012; Bauckhage & Manshaei, 2014). These kernel and multi-layer approaches connect AA with manifold learning techniques. (Javadi & Montanari, 2019) cast NMF as a special case of AA under separability assumptions. In term of algorithmic and scalability improvements  (Chen et al., 2014) proposed a fast active-set algorithm, implemented in the SPAMS toolbox, making AA tractable for large-scale computer vision tasks. (Bauckhage et al., ) adapted a Frank–Wolfe algorithm for AA, reducing runtime. (Mair & Brefeld, 2019) proposed coreset approximations to speed up archetype discovery, and (Abrol & Sharma, 2020) developed a greedy AA algorithm that extended naturally to robust and kernelized variants. Finally, (Mei et al., 2018) proposed an online algorithm for streaming data scenarios.
Mairal et al.’s task-driven dictionary learning framework inspired supervised AA models (Mairal et al., 2011). Related topic modeling approaches such as SurvLDA (Dawson & Kendziorski, 2012) and anchor-based models (Ding et al., 2016; Arora et al., 2013) similarly identify extremal structure. More recently, deep learning has been integrated with AA to handle complex, non-linear data. Wynen et al. (Wynen et al., 2018) applied AA to deep features for archetypal style analysis in artwork. Bauckhage et al. (Bauckhage et al., ) described AA as an autoencoder. Keller et al. (Keller et al., 2020) developed a VAE-based deep AA that accommodates supervision. Applications are widespread. In vision, AA has been used to discover prototypical images (Thurau & Bauckhage, 2009), representative behaviors (Fotiadou et al., 2017), and fMRI brain patterns (Hinrich et al., 2016). In hyperspectral imaging, recent contributions such as SUnAA (Rasti et al., 2023) and entropic descent AA (Zouaoui et al., 2023) set new benchmarks. In NLP, AA was used in multi-document summarization (Canhasi & Kononenko, 2016) and anchor-based topic modeling (Ding et al., 2016; Arora et al., 2013).
Comparison to Our Work
Our method departs from AA by constraining only the decoder to the convex hull of data, while using a linear encoder, maintaining full compatibility with modern SAEs architectures like TopK and JumpReLU. Unlike AA, which jointly optimizes over convex atoms and codes, we retain end to end SAE training. Our goal is not AA approximation or generation, but improved stability and semantic consistency of SAEs. We further introduce a scalable relaxation of the convex constraint, adapted to large-scale training and distinct from prior AA relaxations (e.g., (Mørup & Hansen, 2012)).
Appendix C Formal Definitions of Metrics
@@ C.3 Structure in the Dictionary ($\bm{D}$)
Appendix D Distilling $\bm{A}$ into $\bm{C}$
@@
Appendix E Implementation Details
@@
Appendix F Theoretical properties of Archetypal Dictionary
@@
F.1 Geometric interpretation of A-SAE
Proposition F.1 (Archetypal Dictionary, Convex and Conic Hulls).
@@ F.1 Geometric interpretation of A-SAE
F.2 Stability of Archetypal Dictionary
Proposition F.2 (Geometric Stability of Archetypal Dictionaries).
@@ F.2 Stability of Archetypal Dictionary
F.3 Controlled Rank and Stability in A-SAE Dictionaries